# Bifurcations of the Hénon map with additive bounded noise

### v1

@@ 1. Introduction
Bifurcation theory is pivotal for understanding qualitative changes in dynamical systems. While bifurcation theory is well developed in the context of deterministic dynamical systems, a corresponding theory for random dynamical systems remains relatively unexplored. Existing results on bifurcations in random dynamical systems mainly concern stochastic differential equations with unbounded noise [arnold1995random], even though bounded noise is often realistic from a modelling perspective, for instance in settings where physically relevant fluctuations are intrinsically bounded [d2013bounded].
The aim of this paper is to demonstrate the applicability of a new finite-dimensional map to study bifurcations of attractors of random dynamical systems with bounded noise [kourliouros2023persistence]. We focus on a randomised Hénon map as a prototypical example. Recall that the Hénon map $f:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is given by $f(x,y)=(1-ax^{2}+y,bx)$, depending on parameters $a,b\in\mathbb{R}$ [henon1976two]. We consider this map with additive bounded noise, represented by the random difference equation
@@
contains a neighbourhood of $A$. $A$ is called forward invariant if $z_{i}\in A$ implies $z_{i+1}\in A$ for all noise realisations $\xi_{i}\in\overline{B_{\varepsilon}(0)}$, and it is minimal if no proper subset of $A$ is also forward invariant. Note that this notion of attractor is different from that of so-called random attractors [arnold1995random], which are time-varying objects depending on noise realisations. The attractors we discuss in this paper are deterministic and coincide with the union of fibers of a corresponding random attractor. Crucially, our minimal attractors are interesting from a statistical perspective, since they support ergodic stationary measures [zmarrou2007bifurcations].
Understanding bifurcations of minimal attractors is important beyond the study of random systems. Namely, equations of motion of the form (1.1) also arise in control systems where $(\xi_{i})_{i\in\mathbb{N}_{0}}$ represents a control input. In this context, a minimal attractor is both an invariant control set [colonius2012dynamics] and a reachable set [aubin1984differential, Deimling+1992]. Moreover, bifurcations of minimal attractors are also of relevance to the assessment of resilience of the deterministic system $f$. In particular, [mcgehee1988some, meyer2022intensity] propose the maximal additive perturbation amplitude for which the dynamics remains in the domain of attraction of an attractor $A$, as a measure of resilience. This amplitude coincides with a bifurcation of minimal attractors.
In this paper, we confine the discussion to discrete-time dynamics. The corresponding set-valued dynamics in the continuous time is represented by differential inclusions [aubin1984differential]; see [zmarrou2007bifurcations, homburg2010bifurcations, homburg2013bifurcations] for some results on bifurcations of minimal attractors in that setting.
@@
The compound behaviour of trajectories of the random system (1.1), considering all possible noise realisations, can be analysed through the dynamics of the set-valued map defined by $F(x):=\overline{B_{\varepsilon}(f(x))}$, with the natural extension to sets $S\subset\mathbb{R}^{2}$ as $F(S):=\bigcup_{x\in S}F(x)$. A schematic illustration of this set-valued dynamics is provided in Figure 3. Indeed, minimal attractors of the random system are (minimal) attractors of the set-valued map $F$. This observation has been used to produce Figures 1 and 2, observing the convergence of iterations of initial sets by the set-valued map using set-oriented numerical methods of the software package GAIO [dellnitz2001algorithms]. Unfortunately, however, the approximation of attractors by set-oriented methods is computationally expensive. Moroever, the set-valued point of view does not appear to aid much in the analysis of bifurcations, since set-valued dynamics evolves in the space of compact subsets of $\mathbb{R}^{2}$ which is not a Banach space. This prevents the use of standard bifurcation theory tools in this setting [colonius2012dynamics].
Circumventing the set-valued point of view, an alternative approach to the analysis of boundaries of attractors has been introduced in [kourliouros2023persistence] via a new finite-dimensional boundary map. The main aim of this paper is to provide a proof of concept on how the boundary map can be used to study bifurcations of minimal attractors of random diffeomorphisms with bounded additive noise.
@@ 2. The set-valued map and its bifurcations
If the probability distribution of the noise $(\xi_{i})_{i\in\mathbb{N}_{0}}$ has a non-vanishing Lebesgue density on the $\varepsilon$-ball $\overline{B_{\varepsilon}(0)}$, the support of a stationary measure of (1.1) coincides with a minimal invariant set of the set-valued system $F$ [zmarrou2007bifurcations, araujo2000attractors].
@@
Minimal attractors of the set-valued map can undergo a discontinuous change with respect to the Hausdorff metric, a so-called topological bifurcation [lamb2015topological]*Definition 1.1, as the set-valued map is perturbed. A necessary condition for topological bifurcation is the collision of the minimal invariant set and its dual repeller [lamb2015topological]*Theorem 6.1.
Bifurcation theory in set-valued maps is notoriously challenging [colonius2012dynamics, hans]. Many powerful tools for traditional bifurcation analysis, such as the implicit function theorem, are not readily available for set-valued maps. Given that the collision occurs on the boundary, it is important to approximate and analyse the boundary of the minimal attractor $\partial A$ and its dual repeller $\partial A^{*}$. Studies of the general structure of the attractor’s boundary have proven to be challenging [kourliouros2023persistence, lamb2020boundaries, lamb2021boundaries].
@@ 3. The boundary map
We now consider a set $M$ with smooth boundary $\partial M$. Its unit normal bundle consists of the pairs $(x,n(x))$ of boundary points $x\in\partial M$ and the unit normal $n(x)$ to its tangent space $T_{x}\partial M$. As there are two conventions, of the outer or inner unit normal, we introduce the notation $N_{1}^{\pm}\partial M$ for the corresponding unit normal bundles (where $+$ denotes outer and -- inner).
In the context of this paper, where we consider set-valued maps in the plane $\mathbb{R}^{2}$, the relevant unit normal bundles are isomorphic to $\mathbb{R}^{2}\times S^{1}$. The boundary map $\beta:\mathbb{R}^{2}\times S^{1}\to\mathbb{R}^{2}\times S^{1}$, as introduced in [kourliouros2023persistence, tey2022minimal], is designed to track the outer unit normal bundle of the boundary of a set $M$ under the iterations of the set-valued map $F$ in (2.1).
Definition 3.1 (Boundary Map [kourliouros2023persistence]).
@@
Importantly, the unit normal bundle $N_{1}^{+}\partial A$ of an $F$-invariant set $A$, with continuously differentiable boundary $\partial A$, is $\beta$-invariant [kourliouros2023persistence, tey2022minimal].
@@
Figure 3 illustrates that singularities (points of non-differentiability) can arise on the boundaries of iterates of a set with smooth boundary. Indeed, such singularities can exist also on the boundary $\partial A$ of an invariant set $A$. To characterise the occurrence of such singularities, it is useful to introduce the concept of a contributor [lamb2020boundaries]. We call a point $y\in\partial f(M)$ a contributor of a point $x\in\partial F(M)$, if $\|x-y\|=\varepsilon$.
@@
We now consider the case where the boundary $\partial A$ of a minimal attractor $A$ is continuously differentiable. Due to the uniqueness of contributors of boundary points on $\partial A$, the boundary map $\beta$, when restricted to the normal bundle $\beta|_{N^{+}_{1}\partial A}$ is topologically conjugate to a circle homeomorphism. It is well known that a circle homeomorphism can either consist of periodic points and connections between these periodic points or represent an irrational rotation. In this paper, we focus on the situation where the normal bundle contains a finite number of periodic points of the boundary map $\beta$. It has been established that the normal bundle $N^{+}_{1}\partial A$ is normally attracting [kourliouros2023persistence]. Consequently, the normal bundle includes periodic points of $\beta$ along with the one-dimensional unstable manifolds of saddle periodic points.
@@ 4. Numerical detection of the bifurcations of minimal attractors
In this section, we show that topological and boundary bifurcations of minimal attractors are related to bifurcations of periodic points of the boundary map and their associated one-dimensional invariant manifolds. Topological bifurcations are shown to be related to a fold bifurcation of periodic orbits of the boundary map in Section 4.2, or a heteroclinic bifurcation involving two distinct fixed points in Section 4.3. Analogously for boundary bifurcations, wedge singularities are shown to arise when eigenvalues of periodic orbits change from real-valued to complex-valued in Section 4.1, while the emergence of a cascade of wedge singularities is associated with a non-transversal intersection between the unstable manifold of a saddle fixed point and the strong stable foliations of a stable periodic point, in Section 4.4. The ordering of subsections has been chosen to reflect the increasing complexity of the bifurcation: Section 4.1 and Section 4.2 delve into bifurcations detectable through monitoring the eigenvalues at periodic points, while Section 4.3 and Section 4.4 concerns global bifurcation in the boundary map. The figures presented in this paper can be reproduced using the source codes in [Tey2024]. The set-oriented numerical method is implemented in GAIO [dellnitz2001algorithms] and the approximation of one-dimensional unstable and stable manifolds follows standard techniques (cf. [you1991calculating]).
@@
Figure 7. Boundary of the minimal attractor of the Hénon map with bounded noise (1.1) with parameter values $a=0.13$, $b=0.3$ and $\varepsilon=0.6$, displaying a wedge singularity. On the left, in (a), a numerical approximation of the heteroclinic cycle of the boundary map is depicted, representing the normal bundle of the boundary. The Jacobian of the boundary map at the fixed point $(0.261,-0.455,-2.046)$ on this heteroclinic cycle has eigenvalues $\{0.515,-0.685+0.215i,-0.685-0.215i\}$. These complex eigenvalues cause a spiralling of the unstable manifolds that approach this fixed point in the heteroclinic cycle. The projection of these spiralling manifolds gives rise to a wedge singularity, see the magnified inset in part (a). The projection of the heteroclinic cycle to the state space is depicted in part (b), with the inset showing a magnification of the area with the wedge.
@@
The eigenvalues of the Jacobian at the stable fixed point with coordinates (0.263,0.451,(0.263,-0.451,( 0.263 , - 0.451 , 2.060)-2.060)- 2.060 ) are $\{0.532,-0.768,-0.693\}$. We approach a boundary bifurcation by gradually increasing the value of the parameter $a$ from $a=0.06$ and follow the continuation of the heteroclinic cycle whose projection yields the minimal attractor. The eigenvalues of the Jacobian of the continuation of the fixed point, mentioned above, remain real, until at $a\approx 0.06191$, the two negative eigenvalues collide on the real axis, after which they branch off into the complex plane (as a complex conjugate pair). We observe that the change of stability type of the fixed point, which induces a spiralling behaviour of the connecting unstable manifold to the stable fixed point, creates a wedge singularity on the boundary of the attractor. The situation is sketched in detail in Figure 7 at $a=0.13$.
@@
Figure 8. Sketch of the birth of an isolate wedge singularity in relation to boundary map dynamics. We model the linearised dynamics near a stable fixed point of the boundary map as found in the Hénon map, as in Figure 7, with (a) $a=0.06$ (real eigenvalues) and (b) $a=0.13$ (complex eigenvalues). We consider the linearised dynamics in Jordan normal form, to simplify the presentation: $\left(\begin{smallmatrix}0.532&0&0\\ 0&-0.768&0\\ 0&0&-0.693\end{smallmatrix}\right)$ in (a) and $\left(\begin{smallmatrix}0.515&0&0\\ 0&-0.685&0.215\\ 0&-0.215&-0.685\end{smallmatrix}\right)$ in (b). Blue curves represent smooth invariant manifolds approaching the stable fixed point and their relevant projections are presented in red. The transition from real to complex eigenvalues coincides with the birth of the wedge singularity.
The properties of eigenvalues and eigenvectors mentioned above, give rise to specific properties of the relevant projection. In the case of real eigenvalues, smooth invariant curves approaching the stable fixed point from opposite directions, project to a smooth curve, see Figure 8(a). In the case of complex eigenvalues, the invariant subspace for the Jacobian corresponding to the complex eigenvalues projects to a line. Then, smooth invariant curves spiralling from opposite directions outside of this subspace, give rise to a cascade of intersection points, see Figure 8(b), the first one creating a wedge singularity.
This example shows how a change of stability type of a stable fixed point of the boundary map on the normal bundle of the minimal attractor from real to complex eigenvalues leads to the birth of an isolated wedge singularity. We conjecture that the birth of singularity following a stability type change of the eigenvalues from real to complex generically applies. A mathematical proof, taking into account the relevant geometric setting of the boundary map, requires further formalisation and is beyond the scope of this paper.
@@ 4.2. Topological bifurcation through a fold bifurcation of the boundary map
In Figure 9, we present numerical approximations of the minimal attractors when (a) $a=0.6$ and (b) $a=0.59$. The minimal attractor is depicted in black and its domain of attraction in red. When $a=0.6$, the attractor consists of two disjoint components that are permuted by the dynamics, whereas when $a=0.59$, there is one connected attractor. The transition between these different situations arises discontinuously [lamb2015topological]: when decreasing $a$ from $a=0.6$ to $a=0.59$, around $a=0.595$ the two-component attractor explodes into one connected attractor, in a lower semi-continuous manner [lamb2015topological]. The coloured regions reflect the results of (brute-force) computations using the set-oriented numerical toolbox GAIO [dellnitz2001algorithms], which is well-suited for this task.
@@
We would like to remark that the complexity of the (projection of) relevant unstable manifolds of the boundary map may complicate the understanding of the boundary of the minimal attractor. Indeed, in Figure 9(c) and (d), parts of the projection of the heteroclinic cycle that is deemed irrelevant for the boundary, lying in the interior of the attractor were omitted. In Figure 10, we show the entire projection of the heteroclinic cycle between the two-periodic saddle orbit and the two-periodic stable orbit of the boundary map at $a=0.6$, illustrating the practical complication of deciding which parts of the manifold contribute to the boundary (which can be resolved satisfactorily in a numerical way).
This bifurcation is best appreciated in movie format, tracing superpositions of GAIO data and relevant dynamical objects from the boundary map as $a$ varies. Such movies for this example are provided in [Teyvideo2024] for $a=0.607$ to $a=0.003$, and in [Teyvideoslow2024] for $a=0.607$ to $a=0.58$ in a slower motion, focusing on topological bifurcation.
Figure 10. The projection of the heteroclinic cycle between a saddle two-periodic orbit (red dots) and a stable two-periodic orbit (green dots) of the boundary map for the Hénon map. Only part of the projection corresponds to the boundary of the minimal attractor and the remaining projected manifolds lie in the interior of the minimal attractor. The boundary of the minimal attractor is obtained by removing parts of the projected manifolds in the interior, resulting in the blue curves in Figure 9(c). We note the complexity of the invariant manifolds, complicating the analysis.

### v2

@@ 1. Introduction
Bifurcation theory is pivotal for understanding qualitative changes in dynamical systems. While bifurcation theory is well developed in the context of deterministic dynamical systems, a corresponding theory for random dynamical systems remains relatively unexplored. Existing results on bifurcations in random dynamical systems mainly concern stochastic differential equations with unbounded noise [2], even though bounded noise is often realistic from a modelling perspective, for instance in settings where physically relevant fluctuations are intrinsically bounded [5].
The aim of this paper is to demonstrate the applicability of a new finite-dimensional map to study bifurcations of attractors of random dynamical systems with bounded noise [11]. We focus on a randomised Hénon map as a prototypical example. Recall that the Hénon map $f:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is given by $f(x,y)=(1-ax^{2}+y,bx)$, depending on parameters $a,b\in\mathbb{R}$ [8]. We consider this map with additive bounded noise, represented by the random difference equation
@@
contains a neighbourhood of $A$. AA is called forward invariant if $z_{i}\in A$ implies $z_{i+1}\in A$ for all noise realisations $\xi_{i}\in\overline{B_{\varepsilon}(0)}$, and it is minimal if no proper subset of $A$ is also forward invariant. Note that this notion of attractor is different from that of so-called random attractors [2], which are time-varying objects depending on noise realisations. The attractors we discuss in this paper are deterministic and coincide with the union of fibers of a corresponding random attractor. Crucially, our minimal attractors are interesting from a statistical perspective, since they support ergodic stationary measures [23].
Understanding bifurcations of minimal attractors is important beyond the study of random systems. Namely, equations of motion of the form (1.1) also arise in control systems where $(\xi_{i})_{i\in\mathbb{N}_{0}}$ represents a control input. In this context, a minimal attractor is both an invariant control set [4] and a reachable set [3, 6]. Moreover, bifurcations of minimal attractors are also of relevance to the assessment of resilience of the deterministic system $f$. In particular, [15, 16] propose the maximal additive perturbation amplitude for which the dynamics remains in the domain of attraction of an attractor $A$, as a measure of resilience. This amplitude coincides with a bifurcation of minimal attractors.
In this paper, we confine the discussion to discrete-time dynamics. The corresponding set-valued dynamics in the continuous time is represented by differential inclusions [3]; see [23, 10, 9] for some results on bifurcations of minimal attractors in that setting.
@@
The compound behaviour of trajectories of the random system (1.1), considering all possible noise realisations, can be analysed through the dynamics of the set-valued map defined by $F(x):=\overline{B_{\varepsilon}(f(x))}$, with the natural extension to sets $S\subset\mathbb{R}^{2}$ as $F(S):=\bigcup_{x\in S}F(x)$. A schematic illustration of this set-valued dynamics is provided in Figure 3. Indeed, minimal attractors of the random system are (minimal) attractors of the set-valued map $F$. This observation has been used to produce Figures 1 and 2, observing the convergence of iterations of initial sets by the set-valued map using set-oriented numerical methods of the software package GAIO [7]. Unfortunately, however, the approximation of attractors by set-oriented methods is computationally expensive. Moroever, the set-valued point of view does not appear to aid much in the analysis of bifurcations, since set-valued dynamics evolves in the space of compact subsets of $\mathbb{R}^{2}$ which is not a Banach space. This prevents the use of standard bifurcation theory tools in this setting [4].
Circumventing the set-valued point of view, an alternative approach to the analysis of boundaries of attractors has been introduced in [11] via a new finite-dimensional boundary map. The main aim of this paper is to provide a proof of concept on how the boundary map can be used to study bifurcations of minimal attractors of random diffeomorphisms with bounded additive noise.
@@ 2. The set-valued map and its bifurcations
If the probability distribution of the noise $(\xi_{i})_{i\in\mathbb{N}_{0}}$ has a non-vanishing Lebesgue density on the $\varepsilon$-ball $\overline{B_{\varepsilon}(0)}$, the support of a stationary measure of (1.1) coincides with a minimal invariant set of the set-valued system $F$ [23, 1].
@@
Minimal attractors of the set-valued map can undergo a discontinuous change with respect to the Hausdorff metric, a so-called topological bifurcation [12]*Definition 1.1, as the set-valued map is perturbed. A necessary condition for topological bifurcation is the collision of the minimal invariant set and its dual repeller [12]*Theorem 6.1.
Bifurcation theory in set-valued maps is notoriously challenging [4, 17]. Many powerful tools for traditional bifurcation analysis, such as the implicit function theorem, are not readily available for set-valued maps. Given that the collision occurs on the boundary, it is important to approximate and analyse the boundary of the minimal attractor $\partial A$ and its dual repeller $\partial A^{*}$. Studies of the general structure of the attractor’s boundary have proven to be challenging [11, 13, 14].
@@ 3. The boundary map
We now consider a set $M$ with smooth boundary $\partial M$. Its unit normal bundle consists of the pairs $(x,n(x))$ of boundary points $x\in\partial M$ and the unit normal $n(x)$ to its tangent space $T_{x}\partial M$. As there are two conventions, of the outer or inner unit normal, we introduce the notation $N_{1}^{\pm}\partial M$ for the corresponding unit normal bundles (where ++ denotes outer and - inner).
In the context of this paper, where we consider set-valued maps in the plane $\mathbb{R}^{2}$, the relevant unit normal bundles are isomorphic to $\mathbb{R}^{2}\times S^{1}$. The boundary map $\beta:\mathbb{R}^{2}\times S^{1}\to\mathbb{R}^{2}\times S^{1}$, as introduced in [11, 18], is designed to track the outer unit normal bundle of the boundary of a set $M$ under the iterations of the set-valued map $F$ in (2.1).
Definition 3.1 (Boundary Map [11]).
@@
Importantly, the unit normal bundle $N_{1}^{+}\partial A$ of an $F$-invariant set $A$, with continuously differentiable boundary $\partial A$, is $\beta$-invariant [11, 18].
@@
Figure 3 illustrates that singularities (points of non-differentiability) can arise on the boundaries of iterates of a set with smooth boundary. Indeed, such singularities can exist also on the boundary $\partial A$ of an invariant set $A$. To characterise the occurrence of such singularities, it is useful to introduce the concept of a contributor [13]. We call a point $y\in\partial f(M)$ a contributor of a point $x\in\partial F(M)$, if $\|x-y\|=\varepsilon$.
@@
We now consider the case where the boundary $\partial A$ of a minimal attractor $A$ is continuously differentiable. Due to the uniqueness of contributors of boundary points on $\partial A$, the boundary map $\beta$, when restricted to the normal bundle $\beta|_{N^{+}_{1}\partial A}$ is topologically conjugate to a circle homeomorphism. It is well known that a circle homeomorphism can either consist of periodic points and connections between these periodic points or represent an irrational rotation. In this paper, we focus on the situation where the normal bundle contains a finite number of periodic points of the boundary map $\beta$. It has been established that the normal bundle $N^{+}_{1}\partial A$ is normally attracting [11]. Consequently, the normal bundle includes periodic points of $\beta$ along with the one-dimensional unstable manifolds of saddle periodic points.
@@ 4. Numerical detection of the bifurcations of minimal attractors
In this section, we show that topological and boundary bifurcations of minimal attractors are related to bifurcations of periodic points of the boundary map and their associated one-dimensional invariant manifolds. Topological bifurcations are shown to be related to a fold bifurcation of periodic orbits of the boundary map in Section 4.2, or a heteroclinic bifurcation involving two distinct fixed points in Section 4.3. Analogously for boundary bifurcations, wedge singularities are shown to arise when eigenvalues of periodic orbits change from real-valued to complex-valued in Section 4.1, while the emergence of a cascade of wedge singularities is associated with a non-transversal intersection between the unstable manifold of a saddle fixed point and the strong stable foliations of a stable periodic point, in Section 4.4. The ordering of subsections has been chosen to reflect the increasing complexity of the bifurcation: Section 4.1 and Section 4.2 delve into bifurcations detectable through monitoring the eigenvalues at periodic points, while Section 4.3 and Section 4.4 concerns global bifurcation in the boundary map. The figures presented in this paper can be reproduced using the source codes in [19]. The set-oriented numerical method is implemented in GAIO [7] and the approximation of one-dimensional unstable and stable manifolds follows standard techniques (cf. [22]).
@@
Figure 7. Boundary of the minimal attractor of the Hénon map with bounded noise (1.1) with parameter values $a=0.13$, $b=0.3$ and $\varepsilon=0.6$, displaying a wedge singularity. On the left, in (a), a numerical approximation of the heteroclinic cycle of the boundary map is depicted, representing the normal bundle of the boundary. The Jacobian of the boundary map at the fixed point (0.261,0.455,2.046)(0.261,-0.455,-2.046) on this heteroclinic cycle has eigenvalues $\{0.515,-0.685+0.215i,-0.685-0.215i\}$. These complex eigenvalues cause a spiralling of the unstable manifolds that approach this fixed point in the heteroclinic cycle. The projection of these spiralling manifolds gives rise to a wedge singularity, see the magnified inset in part (a). The projection of the heteroclinic cycle to the state space is depicted in part (b), with the inset showing a magnification of the area with the wedge.
@@
The eigenvalues of the Jacobian at the stable fixed point with coordinates (0.263,0.451,(0.263,-0.451, 2.060)-2.060) are $\{0.532,-0.768,-0.693\}$. We approach a boundary bifurcation by gradually increasing the value of the parameter $a$ from $a=0.06$ and follow the continuation of the heteroclinic cycle whose projection yields the minimal attractor. The eigenvalues of the Jacobian of the continuation of the fixed point, mentioned above, remain real, until at $a\approx 0.06191$, the two negative eigenvalues collide on the real axis, after which they branch off into the complex plane (as a complex conjugate pair). We observe that the change of stability type of the fixed point, which induces a spiralling behaviour of the connecting unstable manifold to the stable fixed point, creates a wedge singularity on the boundary of the attractor. The situation is sketched in detail in Figure 7 at $a=0.13$.
@@
Figure 8. Sketch of the birth of an isolated wedge singularity in relation to boundary map dynamics. We model the linearised dynamics near a stable fixed point of the boundary map as found in the Hénon map, as in Figure 7, with (a) $a=0.06$ (real eigenvalues) and (b) $a=0.13$ (complex eigenvalues). We consider the linearised dynamics in Jordan normal form, to simplify the presentation: $\left(\begin{smallmatrix}0.532&0&0\\ 0&-0.768&0\\ 0&0&-0.693\end{smallmatrix}\right)$ in (a) and $\left(\begin{smallmatrix}0.515&0&0\\ 0&-0.685&0.215\\ 0&-0.215&-0.685\end{smallmatrix}\right)$ in (b). Blue curves represent smooth invariant manifolds approaching the stable fixed point and their relevant projections are presented in red. The transition from real to complex eigenvalues coincides with the birth of the wedge singularity.
The properties of eigenvalues and eigenvectors mentioned above, give rise to specific properties of the relevant projection. In the case of real eigenvalues, smooth invariant curves approaching the stable fixed point from opposite directions, project to a smooth curve, see Figure 8(a). In the case of complex eigenvalues, the invariant subspace for the Jacobian corresponding to the complex eigenvalues projects to a line. Then, smooth invariant curves spiralling toward the fixed point from opposite directions outside of this subspace, give rise to a cascade of transverse intersection points in the projection, the first one creating a wedge singularity as seen in Figure 8(b). The rest of the intersections are too close to the origin, to be visible in this figure.
This example shows how a change of stability type of a stable fixed point of the boundary map on the normal bundle of the minimal attractor from real to complex eigenvalues leads to the birth of an isolated wedge singularity. Indeed, only the first intersection, which creates the wedge, is relevant to the boundary, as the remainder projects to the interior of the attractor, cf. Figure 7(b). We conjecture that the birth of singularity following a stability type change of the eigenvalues from real to complex generically applies. A mathematical proof, taking into account the relevant geometric setting of the boundary map, requires further formalisation and is beyond the scope of this paper.
@@ 4.2. Topological bifurcation through a fold bifurcation of the boundary map
In Figure 9, we present numerical approximations of the minimal attractors when (a) $a=0.6$ and (b) $a=0.59$. The minimal attractor is depicted in black and its domain of attraction in red. When $a=0.6$, the attractor consists of two disjoint components that are permuted by the dynamics, whereas when $a=0.59$, there is one connected attractor. The transition between these different situations arises discontinuously [12]: when decreasing $a$ from $a=0.6$ to $a=0.59$, around $a=0.595$ the two-component attractor explodes into one connected attractor, in a lower semi-continuous manner [12]. The coloured regions reflect the results of (brute-force) computations using the set-oriented numerical toolbox GAIO [7], which is well-suited for this task.
@@
We would like to remark that the complexity of the (projection of) relevant unstable manifolds of the boundary map may complicate the understanding of the boundary of the minimal attractor. Indeed, in Figure 9(c) and (d), parts of the projection of the heteroclinic cycle that is deemed irrelevant for the boundary, lying in the interior of the attractor were omitted. In Figure 10, we show the entire projection of the heteroclinic cycle between the two-periodic saddle orbit and the two-periodic stable orbit of the boundary map at $a=0.6$, illustrating the practical complication of deciding which parts of the manifold contribute to the boundary. While this can be resolved satisfactorily in a numerical way, yielding a good approximation of the boundary, the complexity of the underlying (chaotic) dynamics of the boundary map makes it difficult to obtain the precise structure of the boundary. However, a better understanding of the underlying dynamics may elucidate generic qualitative features.
Figure 10. The projection of the heteroclinic cycle between a saddle two-periodic orbit (red dots) and a stable two-periodic orbit (green dots) of the boundary map for the Hénon map. Only part of the projection corresponds to the boundary of the minimal attractor and the remaining projected manifolds lie in the interior of the minimal attractor. The boundary of the minimal attractor is obtained by removing parts of the projected manifolds in the interior, resulting in the blue curves in Figure 9(c). We observe four wedge singularities on the boundary, two on each disjoint components, see the magnified inset. We note the complexity of the invariant manifolds, complicating the analysis.
This bifurcation is best appreciated in movie format, tracing superpositions of GAIO data and relevant dynamical objects from the boundary map as $a$ varies. Such movies for this example are provided in [20] for $a=0.607$ to $a=0.003$, and in [21] for $a=0.607$ to $a=0.58$ in a slower motion, focusing on topological bifurcation.